# Robotic Learning for Informative Path Planning

### v1

@@ Abstract
AIPP is important to many robotics applications, enabling mobile robots to efficiently collect useful data about initially unknown environments. In addition, learning-based methods are increasingly used in robotics to enhance adaptability, versatility, and robustness across diverse and complex tasks. Our survey explores research on applying robotic learning to adaptive informative path planning (AIPP), bridging the gap between these two research fields. We begin by providing a unified mathematical framework for general AIPP problems. Next, we establish two complementary taxonomies of current work from the perspectives of (i) learning algorithms and (ii) robotic applications. We explore synergies, recent trends, and highlight the benefits of learning-based methods in AIPP frameworks. Finally, we discuss key challenges and promising future directions to enable more generally applicable and robust robotic data-gathering systems through learning. We provide a comprehensive catalogue of papers reviewed in our survey, including publicly available repositories, to facilitate future studies in the field.
@@ 1 Introduction
The field of robotics has witnessed remarkable advancements in recent years, fueled by the growing need for automation and the increasing complexity of tasks required in various application domains. One of the key challenges in robotics is AIPP, which involves planning a trajectory for an autonomous robot to follow that maximizes the information acquired about an unknown environment while simultaneously respecting its resource constraints. This problem is of critical importance in applications such as environmental monitoring, exploration, search and rescue, and inspection across ground, aerial, and aquatic domains (Rayas Fernández et al., 2022; Cao et al., 2023a; Popović et al., 2020; Denniston et al., 2023a; Rückin et al., 2022a; Viseras and Garcia, 2019; Choudhury et al., 2020).
Despite its importance, AIPP is a challenging problem due to the inherent complexity of modeling and predicting new information in the environment, as well as the need to balance the exploration of unknown areas with the exploitation of newly acquired data (Singh et al., 2009). Additionally, noisy sensor measurements introduce uncertainties into the data acquisition process, while actuation uncertainty can lead to deviations from planned trajectories. Moreover, real-world environments are often highly dynamic, which further complicates the modeling process and makes prediction challenging. Decisions must be made in a sequential manner to allow for continual refinement as more information becomes available during a mission.
Conventional approaches such as static pre-computed paths tend to fail on AIPP problems because they often rely on strong assumptions about the environment and cannot adapt to uncertainties or changes online (Galceran and Carreras, 2013). Additionally, these approaches scale poorly to large, complex environments, and may not effectively account for the constraints and capabilities of the robot itself. These conventional solutions applied to the AIPP problem have been limited by computational complexity, lack of adaptability, and an inability to generalize across diverse environments and problems (Tan et al., 2021). The growing popularity of learning-based methods has led to a renewed focus on applying these new techniques to AIPP, offering the promise of more flexible, adaptive, and scalable solutions.
@@
Our review includes different aspects of learning: supervised learning, reinforcement learning, imitation learning, and active learning. We focus on how these techniques can be used for AIPP in robotics. Furthermore, we discuss relevant application domains, such as environmental monitoring, exploration and search, semantic scene understanding, and active simultaneous localization and mapping (SLAM).
While there have been previous surveys addressing aspects of AIPP (Bai et al., 2021; Maboudi et al., 2023; Lluvia et al., 2021; Sung et al., 2023; Taylor et al., 2021; Aniceto and Vivaldini, 2022), our work distinguishes itself by specifically focusing on the intersection of learning-based methods and AIPP. Previous literature reviews have focused on specific subfields of robotic learning. For instance, Taylor et al. (2021) focus on active learning in robotics, highlighting methods suitable for the demands of embodied learning systems. Similarly, Argall et al. (2009) address imitation learning, where robots develop policies from example state-to-action mappings, while Garaffa et al. (2023) investigate reinforcement learning techniques that devise unknown environment exploTration strategies for single and multi-robot exploration. Lauri et al. (2022) exclusively examine the application of Partially Observable Markov Decision Processes (POMDPs) in robotics, and Mukherjee et al. (2022) analyze robotic learning strategies for human-robot collaboration in industrial settings. Other surveys, such as those by Sung et al. (2023), Aniceto and Vivaldini (2022), and Bai et al. (2021) explore different different applications of AIPP, such as environmental monitoring tasks or general path planning techniques (Bai et al., 2021; Galceran and Carreras, 2013; Chen, 2011; Tan et al., 2021; Maboudi et al., 2023). However, these works do not underscore the applicability of learning in adaptive scenarios.
@@
    introduce a standardized mathematical framework for AIPP problems, which provides a unified foundation for understanding and comparing various learning-based methods and application domains;
@@ 2 Mathematical Formulation
We study the general problem of adaptive informative path planning (AIPP). The goal is to find an optimal action sequence $\psi^{*}=(a_{1},\ldots,a_{N})$ of $N$ robot actions $a_{i}\in\mathcal{A},i\in\{1,\ldots,N\},$ where actions may be, e.g., acceleration commands, next robot poses, using a specific sensor, or returning to a charging station. The action sequence $\psi^{*}$ maximizes an information-theoretic criterion $I(\cdot)$:
@@
Our problem setup considers a robot gathering sensor measurements in an initially or partially unknown environment. In this environment $\xi\subset\mathbb{R}^{D}$, AIPP requires online replanning during a mission to adaptively focus on areas of interest as they are discovered. Each point $\mathbf{x}\in\xi$, e.g., a $D$-dimensional pose within the environment, is characterized by its distinctive features, represented as $F(\mathbf{x})\in\mathcal{F}$. Here, $\mathcal{F}$ is a feature space, which could include characteristics like the semantic class, temperature, or radiation level. These features may fluctuate during the mission, as defined by the feature mapping function $F:\xi\to\mathcal{F}$. The notation used to describe our AIPP formulation is summarized in Table 1.
@@
Importantly, the point features $F(\mathbf{x})$, such as temperature, may change over time. Given that the initial environment is unknown, the robot’s understanding of $F(\mathbf{x})$ may need to be updated during the mission. This could result in updating beliefs regarding interesting regions over time as new sensor measurements are received. Thus, to ensure that the robot’s behavior is adaptively informed, we may re-solve the constrained optimization problem in Eq. 1 in a computationally efficient manner during a mission.
@@
$F:\xi\to\mathcal{F}$ feature mapping function semantic class, spatial occupancy, temperature, radiation level
@@ 3.1 Mapping
The AIPP problem assumes the environment $\xi$ to be a priori (partially) unknown. Thus, the true feature mapping $F:\xi\to\mathcal{F}$ and associated true regions of interest are a priori (partially) unknown as well. This makes the original AIPP problem stated in Eq. 1 not only computationally challenging but also impossible to solve directly as exact knowledge about interesting regions requires ground truth access to the feature mapping $F$. Instead of solving Eq. 1 directly, we model a stochastic process $\tilde{F}$ over all possible feature mapping functions. At each mission time step $t$, the robot’s belief about the true feature mapping $F$, i.e., the probability measure associated with $\tilde{F}$ conditioned on the action-observation history, is updated by:
@@
To approximate the AIPP problem in Eq. 1 as closely as possible, the robot’s belief about $\tilde{F}$ in Eq. 2 may be updated online during a mission as new sensor data $z_{t}$ arrives at time step $t$ after executing a sensing action $a_{t}$. Note that the measurements $z_{t}$ at time step $t$ may be uncertain or noisy. The measurements $z_{t}$ are assumed to be sampled from $p(z\,|\,F)$ representing a probabilistic sensor model. Furthermore, there may be uncertainty in localization and actuation, leading to imperfect estimation of the robot pose as measurements are taken. In Sec. 4 and Sec. 5, we discuss in detail how these different sources of uncertainty can be accounted for in the context of learning-based methods.
To compute the belief $\tilde{F}$ following Eq. 2 based on the collection of stochastic measurements $\{z_{1},\ldots,z_{t}$}, various methods for environment mapping are commonly used. Table 2 provides an overview of commonly used mapping methods and corresponding metrics derived from them used to evaluate the performance of AIPP methods. We consider four broad categories in our taxonomy: occupancy grid mapping (Elfes, 1989), Gaussian processes (Rasmussen and Williams, 2006), graph-based methods, and implicit neural representations (Mildenhall et al., 2021). In general, the chosen mapping method for a given AIPP scenario depends on the environment, task, and available computing resources. For instance, Gaussian processes are applicable for problems where the to-be-mapped features $F(\mathbf{x})$ support continuous feature spaces $\mathcal{F}$ and the map supports continuous environment representations $\xi$, whereas occupancy grid mapping is suitable for discrete mapping tasks. In contrast, graph-based methods are typically used for active SLAM problems (Placed et al., 2023) where both environment mapping and robot localization are concurrent. Recently, implicit neural representations, such as NeRFs (Mildenhall et al., 2021) and pixelNeRFs (Yu et al., 2021), are gaining popularity for AIPP without relying on an explicit environment map. Due to their lightweight nature, these methods enable conserving computational and memory resources during planning in 3D reconstruction tasks where a high level of detail is required.
@@
Occupancy grid Entropy (Popović et al., 2020; Westheider et al., 2023; Bai et al., 2017; Chen et al., 2020; Sukkar et al., 2019; Choudhury et al., 2018)
Mutual information (Lodel et al., 2022; Chen et al., 2019)
Root mean squared error (Popović et al., 2020)
Mean absolute error (Chen et al., 2021; Viseras et al., 2021)
Mean intersection-over-union (Georgakis et al., 2022a; Ramakrishnan et al., 2020)
F1-score (Westheider et al., 2023)
Accuracy (Schmid et al., 2023; Georgakis et al., 2022a; Ramakrishnan et al., 2020)
Precision/recall (Schmid et al., 2023)
Coverage (Schmid et al., 2023, 2022; Niroui et al., 2019; Cao et al., 2023b; Reinhart et al., 2020; Zeng et al., 2022; Shrestha et al., 2019; Chen et al., 2021; Zacchini et al., 2023; Liu et al., 2023; Vasquez-Gomez et al., 2021; Mendoza et al., 2020; Caley et al., 2019; Georgakis et al., 2022a; Ramakrishnan et al., 2020; Chen et al., 2022; Tao et al., 2023; Li et al., 2023; Gao and Zhang, 2022; Sukkar et al., 2019; Zwecher et al., 2022; Dai et al., 2022)
Observed surfaces (Hepp et al., 2018; Dhami et al., 2023; Pan et al., 2023a)
Landmark error (Choudhury et al., 2020)
Gaussian process Entropy (Denniston et al., 2023b)
Expected improvement (Denniston et al., 2023b; Song et al., 2015)
Probability of improvement (Denniston et al., 2023b)
Covariance matrix trace (Cao et al., 2023a; Rückin et al., 2022a; Popović et al., 2020, 2020; Ott et al., 2023; Yang et al., 2023a; Yanes Luis et al., 2023)
Mutual information (Hitz et al., 2017; Wei and Zheng, 2020)
Root mean squared error (Cao et al., 2023a; Rückin et al., 2022a; Rayas Fernández et al., 2022; Vivaldini et al., 2019; Popović et al., 2020, 2020; Hollinger and Sukhatme, 2014; Viseras and Garcia, 2019; Yanes Luis et al., 2023; Choi and Cielniak, 2021)
Mean log likelihood (Popović et al., 2020)
Distribution divergence (Cao et al., 2023c)
Mean uncertainty (Cao et al., 2023c; Hollinger and Sukhatme, 2014)
Upper confidence bound (Marchant et al., 2014)
Graph-based Landmark uncertainty (Chen et al., 2021, 2020)
Landmark error (Chen et al., 2020)
Localization uncertainty (Chen et al., 2020; Popović et al., 2020)
Localization error (Popović et al., 2020; Hüttenrauch et al., 2019)
Target uncertainty (Duecker et al., 2021; Tzes et al., 2023; Singh et al., 2009)
Detection rate (Best et al., 2019)
Path costs (Meliou et al., 2007)
Neural representation Peak signal-to-noise ratio (Jin et al., 2023; Pan et al., 2022; Ran et al., 2022; Zhan et al., 2023; Sünderhauf et al., 2023; Pan et al., 2023b)
Structural similarity index measure (Jin et al., 2023; Pan et al., 2022; Ran et al., 2022; Zhan et al., 2023; Pan et al., 2023b)
Learned perceptual image patch similarity (Pan et al., 2022; Ran et al., 2022; Zhan et al., 2023)
Accuracy (Ran et al., 2022; Zhan et al., 2023; Yan et al., 2023)
Completion (Ran et al., 2022; Zhan et al., 2023; Yan et al., 2023)
Completion ratio (Ran et al., 2022; Yan et al., 2023)
F1-score (Zhan et al., 2023; Lee et al., 2022)
Chamfer distance (Zhan et al., 2023)
Precision (Lee et al., 2022)
Recall (Lee et al., 2022)
Surface coverage (Lee et al., 2022; Yan et al., 2023)
@@ 3.2 Evaluation Metrics
The goal of AIPP is to adaptively acquire information about initially unknown environment features $F$ by collecting uncertain or noisy measurements $z_{t}$ during a mission, usually captured as a posterior map belief $\tilde{F}$ in Eq. 2. Because the robot mission goals are task- and application-dependent, e.g., searching for targets (Ott et al., 2023; Tzes et al., 2023; Cao et al., 2023c) or regions of high surface temperature (Rückin et al., 2022a; Meliou et al., 2007; Westheider et al., 2023), there is a lack of standard evaluation metrics used to quantify the performance of AIPP algorithms. Most works choose evaluation metrics closely related to their specified information criterion $I$ in Eq. 1 because they reflect or approximate the mission goal. Table 2 lists commonly used evaluation metrics. Although these metrics vary significantly in the AIPP community, most works consider quantitative measures based on the map representation to compute the belief $\tilde{F}$, as reflected by our taxonomy.
@@
Applications aiming to collect information about discrete features $F$, such as spatial occupancy (Bai et al., 2017; Chen et al., 2020) or land cover classification Popović et al. (2020); Vivaldini et al. (2019); Rayas Fernández et al. (2022), commonly use the entropy of the belief $\tilde{F}$ as a measure for the remaining uncertainty of the environment features (Popović et al., 2020; Westheider et al., 2023; Bai et al., 2017; Chen et al., 2020; Sukkar et al., 2019; Choudhury et al., 2018). Lower entropy of $\tilde{F}$ indicates better AIPP performance. Following popular computer vision benchmarks Everingham et al. (2010); Cordts et al. (2016), some works evaluate the classification quality of $\tilde{F}$ by computing the mean Intersection-over-Union (mIoU(Georgakis et al., 2022a; Ramakrishnan et al., 2020), accuracy (Schmid et al., 2023; Georgakis et al., 2022a; Ramakrishnan et al., 2020), or F1-score (Westheider et al., 2023) of the maximum a posteriori estimate of the map belief $\tilde{F}$ given the ground truth features $F$.
@@
Methods gathering information about continuous features $F$ in an environment, e.g., bacteria level in a lake (Hitz et al., 2017) or magnetic field strength (Viseras and Garcia, 2019), often use Gaussian process models Hitz et al. (2017); Cao et al. (2023a); Viseras and Garcia (2019); Ott et al. (2023) or some variant of Bayesian filtering Popović et al. (2020); Rückin et al. (2022a); Choudhury et al. (2020); Denniston et al. (2023a) as the map representations. Although Gaussian processes allow for computing the differential entropy of $\tilde{F}$ in closed-form (Williams and Rasmussen, 2006), it is common to use the covariance matrix trace to approximate the remaining uncertainty about the environment features and hence the information criterion $I$ due to its computational efficiency. Despite relaxed computational requirements, most works evaluate AIPP performance based on the covariance matrix trace during offline evaluation as well (Cao et al., 2023a; Rückin et al., 2022a; Popović et al., 2020, 2020; Ott et al., 2023; Yang et al., 2023a; Yanes Luis et al., 2023). To evaluate the reconstruction quality of $\tilde{F}$, the root mean squared error (RMSE) between the maximum a posteriori estimate of $\tilde{F}$ and the ground truth $F$ is usually reported Cao et al. (2023a); Rückin et al. (2022a); Rayas Fernández et al. (2022); Vivaldini et al. (2019); Popović et al. (2020, 2020); Hollinger and Sukhatme (2014); Viseras and Garcia (2019); Yanes Luis et al. (2023); Choi and Cielniak (2021).
@@
Evaluation methods for implicit neural representations differ in that they only rely on a trained neural network to represent the entire environment. The metrics can be grouped based on two aspects: the quality of the rendered images and the quality of the geometry of the reconstructed surface. Image quality metrics, e.g., peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM), and learned perceptual image patch similarity (LPIPS), are computed based on a set of evenly distributed test images in the environment (Jin et al., 2023; Pan et al., 2022; Sünderhauf et al., 2023; Ran et al., 2022; Zhan et al., 2023; Pan et al., 2023b). To evaluate the reconstruction quality, the learned geometric representation is usually first post-processed, e.g., by using the marching cubes algorithm to extract a mesh. Then, a set of points is sampled on the mesh to compute various metrics, e.g., mesh accuracy or completion (Lee et al., 2022; Ran et al., 2022; Zhan et al., 2023).
@@
Some methods do not assume the robot to have perfect ground truth localization information during missions but additionally account for localization uncertainty while building the map belief $\tilde{F}$. These works propagate localization uncertainty into the map belief update in Eq. 2 using graph-based SLAM systems (Chen et al., 2020, 2021; Popović et al., 2020; Tzes et al., 2023) and Gaussian process variants supporting input uncertainty (Popović et al., 2020). In addition to assessing the quality of the map, the quality of robot localization is also evaluated to determine whether the AIPP method can gather not only informative measurements to improve the environment map, but also measurements that improve localization to guarantee the construction of an accurate map. Some works compute the trajectory errors between the estimated and ground truth executed trajectory Chen et al. (2021). Other works track the localization uncertainty of the robot pose belief Popović et al. (2020); Chen et al. (2020), the localization error between the estimated and ground truth robot pose Popović et al. (2020); Hüttenrauch et al. (2019), or the landmark errors between estimated and ground truth landmark poses (Chen et al., 2020).
@@
An alternative problem formulation considers using AIPP to actively improve the sensor model reasoning about measurements $z_{t}$. In Eq. 1, this is done by linking the information criterion $I$ to the value of new training data acquired during a mission. Such formulations generally do not aim to improve the belief of the map $\tilde{F}$ representing a physically quantifiable environmental phenomenon. Instead, works in this category target improving a neural network for semantic segmentation using RGB images (Blum et al., 2019; Zurbrügg et al., 2022; Rückin et al., 2023, 2024) or object detection (Chaplot et al., 2021). AIPP performance is evaluated using a held-out test set to quantify the neural network prediction performance Everingham et al. (2010); Cordts et al. (2016). Commonly used metrics include mIoU Blum et al. (2019); Zurbrügg et al. (2022); Rückin et al. (2023), F1-score (Rückin et al., 2023, 2024), or average precision (Chaplot et al., 2021), depending on the prediction task.
@@
Due to the sequential nature of the AIPP problem, most works consider a well-performing AIPP approach not only to show strong evaluation metrics after a mission is over but also to ensure fast improvement during a mission as the remaining mission budget $B$ decreases. Trends in performance over mission time allow us to judge the efficiency of budget allocation, which is critical in AIPP as reflected by the constraints in Eq. 1. Common evaluation strategies include analyzing how the performance-based evaluation metrics evolve with the remaining mission budget, e.g., trajectory steps (Ott et al., 2023; Westheider et al., 2023; Viseras and Garcia, 2019; Cao et al., 2023a; Denniston et al., 2023a) or energy-based budgets (Popović et al., 2020; Rückin et al., 2022a; Hollinger and Sukhatme, 2014; Choudhury et al., 2020).

### v3

# Learning-based Methods for Adaptive Informative Path Planning

@@ Abstract
AIPP is important to many robotics applications, enabling mobile robots to efficiently collect useful data about initially unknown environments. In addition, learning-based methods are increasingly used in robotics to enhance adaptability, versatility, and robustness across diverse and complex tasks. Our survey explores research on applying robotic learning to adaptive informative path planning (AIPP), bridging the gap between these two research fields. We begin by providing a unified mathematical problem definition for general AIPP problems. Next, we establish two complementary taxonomies of current work from the perspectives of (i) learning algorithms and (ii) robotic applications. We explore synergies, recent trends, and highlight the benefits of learning-based methods in AIPP frameworks. Finally, we discuss key challenges and promising future directions to enable more generally applicable and robust robotic data-gathering systems through learning. We provide a comprehensive catalog of papers reviewed in our survey, including publicly available repositories, to facilitate future studies in the field.
@@ 1 Introduction
The field of robotics has witnessed remarkable advancements in recent years, fueled by the growing need for automation and the increasing complexity of tasks required in various application domains. One of the key challenges in robotics is AIPP, which involves planning a trajectory for an autonomous robot to follow that maximizes the information acquired about an unknown environment while simultaneously respecting its resource constraints. This problem is of critical importance in applications such as environmental monitoring, exploration, search and rescue, and inspection across ground, aerial, and aquatic domains [1, 2, 3, 4, 5, 6, 7].
Despite its importance, AIPP is a challenging problem due to the inherent complexity of modeling and predicting new information in the environment, as well as the need to balance the exploration of unknown areas with the exploitation of newly acquired data [8]. Additionally, noisy sensor measurements introduce uncertainties into the data acquisition process, while actuation uncertainty can lead to deviations from planned trajectories. Moreover, real-world environments are often highly dynamic, which further complicates the modeling process and makes prediction challenging. Decisions must be made in a sequential manner to allow for continual refinement as more information becomes available during a mission.
Conventional approaches such as static pre-computed paths tend to fail on AIPP problems because they often rely on strong assumptions about the environment and cannot adapt to uncertainties or changes online [9]. Additionally, these approaches scale poorly to large, complex environments, and may not effectively account for the constraints and capabilities of the robot itself. These conventional solutions applied to the AIPP problem have been limited by computational complexity, lack of adaptability, and an inability to generalize across diverse environments and problems [10]. The growing popularity of learning-based methods has led to a renewed focus on applying these new techniques to AIPP, offering the promise of more flexible, adaptive, and scalable solutions.
Figure 1: Overview of our survey paper. We review related work in learning-based methods for adaptive informative path planning (AIPP) from the perspectives of learning-based approaches (left) and practical applications (right).
@@
Fig. 1 shows an overview of the main topics addressed in our survey paper. Our review includes different aspects of learning: supervised learning, reinforcement learning, imitation learning, and active learning. We focus on how these techniques can be used for AIPP in robotics. Furthermore, we discuss relevant application domains, such as environmental monitoring, exploration and search, semantic scene understanding, and active simultaneous localization and mapping (SLAM).
While there have been previous surveys addressing aspects of AIPP [11, 12, 13, 14, 15, 16], our work distinguishes itself by specifically focusing on the intersection of learning-based methods and AIPP. Previous literature reviews have focused on specific subfields of robotic learning. For instance, Taylor et al. [15] focus on active learning in robotics, highlighting methods suitable for the demands of embodied learning systems. Similarly, Argall et al. [17] address imitation learning, where robots develop policies from example state-to-action mappings, while Garaffa et al. [18] investigate reinforcement learning techniques that devise unknown environment exploration strategies for single and multi-robot exploration. Lauri et al. [19] exclusively examine the application of Partially Observable Markov Decision Processes (POMDPs) in robotics, and Mukherjee et al. [20] analyze robotic learning strategies for human-robot collaboration in industrial settings. Other surveys, such as those by Sung et al. [14], Aniceto and Vivaldini [16], and Bai et al. [11] explore different different applications of AIPP, such as environmental monitoring tasks or general path planning techniques [11, 9, 21, 10, 12]. However, these works do not underscore the applicability of learning in adaptive scenarios.
@@
    introduce a standardized mathematical problem definition for AIPP tasks, which provides a unified foundation for understanding and comparing various learning-based methods and application domains;
@@ 2 Mathematical Formulation
First, we develop an underlying mathematical formulation to encompass the works included in our survey. We study the general problem of adaptive informative path planning (AIPP). The goal is to find an optimal action sequence $\psi^{*}=(a_{1},\ldots,a_{N})$ of $N$ robot actions $a_{i}\in\mathcal{A},i\in\{1,\ldots,N\},$ where actions may be, e.g., acceleration commands, next robot poses, using a specific sensor, or returning to a charging station. The action sequence $\psi^{*}$ maximizes an information-theoretic criterion $I(\cdot)$:
@@
Our problem setup considers a robot gathering sensor measurements in an initially or partially unknown environment. In this environment $\xi\subset\mathbb{R}^{D}$, AIPP requires online replanning during a mission to adaptively focus on areas of interest as they are discovered. Each point $\mathbf{x}\in\xi$, e.g., a $D$-dimensional pose within the environment, is characterized by its distinctive features, represented as $F(\mathbf{x},t)\in\mathcal{F}$. Here, $\mathcal{F}$ is a feature space, which could include characteristics like the semantic class, temperature, or radiation level, and may vary over time. These features may fluctuate during the mission, as defined by the feature mapping function $F:\xi\times\mathbb{R}\to\mathcal{F}$. The notation used to describe our AIPP formulation is summarized in Table 1.
@@
Importantly, the point features $F(\mathbf{x},t)$, such as temperature, may change over time. Given that the initial environment is unknown, the robot’s understanding of $F(\mathbf{x},t)$ may need to be updated during the mission. This could result in updating beliefs regarding interesting regions over time as new sensor measurements are received. Thus, to ensure that the robot’s behavior is adaptively informed, we may re-solve the constrained optimization problem in Eq. 1 in a computationally efficient manner during a mission.
@@
$F:\xi\times\mathbb{R}\to\mathcal{F}$ feature mapping function semantic class, spatial occupancy, temperature, radiation level
@@ 3.1 Mapping
The AIPP problem assumes the environment $\xi$ to be a priori (partially) unknown. Thus, the true feature mapping $F:\xi\times\mathbb{R}\to\mathcal{F}$ and associated true regions of interest are a priori (partially) unknown as well. This makes the original AIPP problem stated in Eq. 1 not only computationally challenging but also impossible to solve directly as exact knowledge about interesting regions requires ground truth access to the feature mapping $F$. Instead of solving Eq. 1 directly, we model a stochastic process $\tilde{F}$ over all possible feature mapping functions. At each mission time step $t$, the robot’s belief about the true feature mapping $F$, i.e., the probability measure associated with $\tilde{F}$ conditioned on the action-observation history, is updated by:
@@
The robot’s belief about $\tilde{F}$ in Eq. 2 may be updated online during a mission as new sensor data $z_{t}$ arrives at time step $t$ after executing a sensing action $a_{t}$. Note that the information criterion $I(\cdot)$ reflects the planning objective and is computed based on the map state. In general, the computation of certain information criteria, e.g. entropy, may change for different map representations, e.g. occupancy maps and Gaussian processes, but the key underlying theoretical concepts remain the same.
The measurements $z_{t}$ obtained at time step $t$ may be uncertain or noisy. They are assumed to be sampled from $p(z\,|\,F)$ representing a probabilistic sensor model. Furthermore, there may be uncertainty in localization and actuation, leading to imperfect estimation of the robot pose as measurements are taken. In Sec. 4 and Sec. 5, we discuss in detail how these different sources of uncertainty can be accounted for in the context of learning-based methods.
To compute the belief $\tilde{F}$ following Eq. 2 based on the collection of stochastic measurements $\{z_{1},\ldots,z_{t}$}, various methods for environment mapping are commonly used. Table 2 provides an overview of commonly used mapping methods and corresponding metrics derived from them used to evaluate the performance of AIPP methods. We consider four broad categories in our taxonomy: occupancy grid mapping [22], Gaussian processes [23], graph-based methods, and implicit neural representations [24]. In general, the chosen mapping method for a given AIPP scenario depends on the environment, task, and available computing resources. For instance, Gaussian processes are applicable for problems where the to-be-mapped features $F(\mathbf{x},t)$ support continuous feature spaces $\mathcal{F}$ and the map supports continuous environment representations $\xi$, whereas occupancy grid mapping is suitable for discrete mapping tasks. In contrast, graph-based methods are typically used for active SLAM problems [25] where both environment mapping and robot localization are concurrent. Recently, implicit neural representations, such as NeRFs [24] and pixelNeRFs [26], are gaining popularity for AIPP without relying on an explicit environment map. Due to their lightweight nature, these methods enable conserving computational and memory resources during planning in 3D reconstruction tasks where a high level of detail is required.
@@
Occupancy grid Entropy [3, 27, 28, 29, 30, 31]
Mutual information [32, 33]
Root mean squared error [3]
Mean absolute error [34, 35]
Mean intersection-over-union [36, 37]
F1-score [27]
Accuracy [38, 36, 37]
Precision/recall [38]
Coverage [38, 39, 40, 41, 42, 43, 44, 34, 45, 46, 47, 48, 49, 36, 37, 50, 51, 52, 53, 30, 54, 55]
Observed surfaces [56, 57, 58]
Landmark error [7]
Gaussian process Entropy [59]
Expected improvement [59, 60]
Probability of improvement [59]
Covariance matrix trace [2, 5, 3, 61, 62, 63, 64]
Mutual information [65, 66]
Root mean squared error [2, 5, 1, 67, 3, 61, 68, 6, 64, 69]
Mean log likelihood [3]
Distribution divergence [70]
Mean uncertainty [70, 68]
Upper confidence bound [71]
Graph-based Landmark uncertainty [34, 29]
Landmark error [29]
Localization uncertainty [29, 61]
Localization error [61, 72]
Target uncertainty [73, 74, 8]
Detection rate [75]
Path costs [76]
Neural representation Peak signal-to-noise ratio [77, 78, 79, 80, 81, 82]
Structural similarity index measure [77, 78, 79, 80, 82]
Learned perceptual image patch similarity [78, 79, 80]
Accuracy [79, 80, 83]
Completion [79, 80, 83]
Completion ratio [79, 83]
F1-score [80, 84]
Chamfer distance [80]
Precision [84]
Recall [84]
Surface coverage [84, 83]
@@ 3.2 Evaluation Metrics
The goal of AIPP is to adaptively acquire information about initially unknown environment features $F$ by collecting uncertain or noisy measurements $z_{t}$ during a mission, usually captured as a posterior map belief $\tilde{F}$ in Eq. 2. Because the robot mission goals are task- and application-dependent, e.g., searching for targets [62, 74, 70] or regions of high surface temperature [5, 76, 27], there is a lack of standard evaluation metrics used to quantify the performance of AIPP algorithms. Most works choose evaluation metrics closely related to their specified information criterion $I$ in Eq. 1 because they reflect or approximate the mission goal. Table 2 lists commonly used evaluation metrics. Although these metrics vary significantly in the AIPP community, most works consider quantitative measures based on the map representation to compute the belief $\tilde{F}$, as reflected by our taxonomy.
@@
Applications aiming to collect information about discrete features $F$, such as spatial occupancy [28, 29] or land cover classification [3, 67, 1], commonly use the entropy of the belief $\tilde{F}$ as a measure for the remaining uncertainty of the environment features [3, 27, 28, 29, 30, 31]. Lower entropy of $\tilde{F}$ indicates better AIPP performance. Following popular computer vision benchmarks [85, 86], some works evaluate the classification quality of $\tilde{F}$ by computing the mean Intersection-over-Union (mIoU[36, 37], accuracy [38, 36, 37], or F1-score [27] of the maximum a posteriori estimate of the map belief $\tilde{F}$ given the ground truth features $F$.
@@
Methods gathering information about continuous features $F$ in an environment, e.g., bacteria level in a lake [65] or magnetic field strength [6], often use Gaussian process models [65, 2, 6, 62] or some variant of Bayesian filtering [3, 5, 7, 4] as the map representations. Although Gaussian processes allow for computing the differential entropy of $\tilde{F}$ in closed-form [87], it is common to use the covariance matrix trace to approximate the remaining uncertainty about the environment features and hence the information criterion $I$ due to its computational efficiency. Despite relaxed computational requirements, most works evaluate AIPP performance based on the covariance matrix trace during offline evaluation as well [2, 5, 3, 61, 62, 63, 64]. To evaluate the reconstruction quality of $\tilde{F}$, the root mean squared error (RMSE) between the maximum a posteriori estimate of $\tilde{F}$ and the ground truth $F$ is usually reported [2, 5, 1, 67, 3, 61, 68, 6, 64, 69].
@@
Evaluation methods for implicit neural representations differ in that they only rely on a trained neural network to represent the entire environment. The metrics can be grouped based on two aspects: the quality of the rendered images and the quality of the geometry of the reconstructed surface. Image quality metrics, e.g., peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM), and learned perceptual image patch similarity (LPIPS), are computed based on a set of evenly distributed test images in the environment [77, 78, 81, 79, 80, 82]. To evaluate the reconstruction quality, the learned geometric representation is usually first post-processed, e.g., by using the marching cubes algorithm to extract a mesh. Then, a set of points is sampled on the mesh to compute various metrics, e.g., mesh accuracy or completion [84, 79, 80].
@@
Some methods do not assume the robot to have perfect ground truth localization information during missions but additionally account for localization uncertainty while building the map belief $\tilde{F}$. These works propagate localization uncertainty into the map belief update in Eq. 2 using graph-based SLAM systems [29, 34, 61, 74] and Gaussian process variants supporting input uncertainty [61]. In addition to assessing the quality of the map, the quality of robot localization is also evaluated to determine whether the AIPP method can gather not only informative measurements to improve the environment map, but also measurements that improve localization to guarantee the construction of an accurate map. Some works compute the trajectory errors between the estimated and ground truth executed trajectory [34]. Other works track the localization uncertainty of the robot pose belief [61, 29], the localization error between the estimated and ground truth robot pose [61, 72], or the landmark errors between estimated and ground truth landmark poses [29].
@@
An alternative problem formulation considers using AIPP to actively improve the sensor model reasoning about measurements $z_{t}$. In Eq. 1, this is done by linking the information criterion $I$ to the value of new training data acquired during a mission. Such formulations generally do not aim to improve the belief of the map $\tilde{F}$ representing a physically quantifiable environmental phenomenon. Instead, works in this category target improving a neural network for semantic segmentation using RGB images [88, 89, 90, 91] or object detection [92]. AIPP performance is evaluated using a held-out test set to quantify the neural network prediction performance [85, 86]. Commonly used metrics include mIoU [88, 89, 90], F1-score [90, 91], or average precision [92], depending on the prediction task.
@@
Due to the sequential nature of the AIPP problem, most works consider a well-performing AIPP approach not only to show strong evaluation metrics after a mission is over but also to ensure fast improvement during a mission as the remaining mission budget $B$ decreases. Trends in performance over mission time allow us to judge the efficiency of budget allocation, which is critical in AIPP as reflected by the constraints in Eq. 1. Common evaluation strategies include analyzing how the performance-based evaluation metrics evolve with the remaining mission budget, e.g., trajectory steps [62, 27, 6, 2, 4] or energy-based budgets [3, 5, 68, 7].